OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = x^3 + 4 and g(x) = cube root(x - 4)

OpenStudy (mathstudent55):

\(f(x) = x^3 + 4 \) \(g(x) = \sqrt[3]{x - 4} \) \(f(g(x)) = f(\sqrt[3]{x - 4}) \) \( ~~~~~~~~~~~~~= ( \sqrt[3]{x - 4})^3 + 4 \) \(~~~~~~~~~~~~~= x - 4 + 4\) \(~~~~~~~~~~~~~ = x\) Now try doing the same for \(g(f(x))\).

OpenStudy (anonymous):

your job is to compute \[f(g(x))=f(\sqrt[3]{x-4})=(\sqrt[3]{x-4})^3+4\] and see that you get \(x\)

OpenStudy (anonymous):

oh, what @mathstudent55 said!